OpenStudy (anonymous):

The Hansons have reduced the average number of hours they watch television each day from 5.5 hours in March to 2.5 hours in June. Find the rate of change in the average number of hours they watch television each day.

ganeshie8 (ganeshie8):

change in tv watching hours = \(2.5 - 5.5 = -3\)

ganeshie8 (ganeshie8):

there were \(3 \) months in between

ganeshie8 (ganeshie8):

\(\textbf{rate of change of something = }\frac{\textbf{change in something}}{time}\)

ganeshie8 (ganeshie8):

\(\huge = \frac{-3}{3} \\=-1\)

ganeshie8 (ganeshie8):

this just means this : the hansons are reducing their tv watching hours at the rate of "-1 hour" per month. hope it is clear..